There are 24 students in a class. Each student is wearing either a hat or gym shoes or both.
16 are wearing hats
20 are wearing shoes
How many are wearing both?

Answer by Edwin McCravy(20054)

There are 24 students in a class. Each student is wearing either a hat or gym shoes or both.
16 are wearing hats
20 are wearing shoes
How many are wearing both?

Use the formula:  N(A or B) = N(A) + N(B) - N(A and B)
Where N() means the number of 

A is the event of wearing a hat.
B is the event of wearing gym shoes.

 N(A or B) = N(A) + N(B) - N(A and B)
        24 =  16  +  20  - N(A and B)
        24 =  36 - N(A and B)
N(A and B) = 12
 
Edwin